OpenStudy (anonymous):

Solve the equation by completing the square: x^2 + 2x - 8 = 0 A. x = 2 and x = -8 B. x = -4 and x = 2 C. x = 8 D. x = 1, x = 2 and x = -8

OpenStudy (anonymous):

x^2+4x-2x-8=0 x(x+4)-2(x+4)=0 =>x-2=0 or x+4=0 x=2 or x=-4 answer is B